# Mutual Information Estimation via Score-to-Fisher Bridge for Nonlinear Gaussian Noise Channels

### v2

@@ Abstract
We present a numerical method to evaluate mutual information (MI) in nonlinear Gaussian noise channels by using denoising score matching (DSM) learning for estimating the score function of channel output. Via de Bruijn’s identity, Fisher information estimated from the learned score function yields accurate estimates of MI through a Fisher integral representation for a variety of priors and channel nonlinearities. In this work, we propose a comprehensive theoretical foundation for the Score-to-Fisher bridge methodology, along with practical guidelines for its implementation. We also conduct extensive validation experiments, comparing our approach with closed-form solutions and a kernel density estimation baseline. The results of our numerical experiments demonstrate that the proposed method is both practical and efficient for MI estimation in nonlinear Gaussian noise channels.
@@ I Introduction
$Y_{t}=f(X)+Z_{t},\qquad Z_{t}\sim\mathcal{N}(\mathbf{0},t\mathbf{I}_{n}),$ (1)
@@
so that the mutual information $I(X;Y_{t})$ can be obtained by integrating the Fisher information $J(Y_{t})$.
@@
In this paper, we propose a score-based estimator of MI tailored to $Y_{t}=f(X)+Z_{t}$ [^1] A software implementation to reproduce the numerical results in this paper is available at https://github.com/wadayama/score-to-fisher-mi. The method learns the score of channel output, or the marginal density $p_{Y_{t}}$, using DSM learning and then obtain an estimate of the Fisher information $\widehat{J}(Y_{t}).$ From this, we obtain an estimate of the mutual information via an Fisher integral representation (4). Crucially, the approach never requires access to the posterior $p_{X|Y}(\cdot|\cdot)$. All we need is possibility of forward channel simulation, i.e., we assume that generation of samples $\bm{x}\sim p_{X}(\cdot)$, evaluations of $\bm{w}=f(\bm{x})$, and Gaussian noise addition are computationally feasible.
The main contributions of this paper are summarized as follows:
  • Score-only MI estimation for nonlinear front-ends: We formalize a pipeline— DSM learning of $p_{Y_{t}}$ $\Rightarrow$ Fisher estimation $\Rightarrow$ Fisher integral representation—that applies to arbitrary deterministic $f$ as soon as $p_{Y_{t}}$ can be easily sampled.
  • Practical recipe: We detail implementation choices that make the estimator stable and reproducible: logarithmic grids in $t$, tail corrections for the MI integral.
  • Validation and use-cases: We verify the estimator on Gaussian inputs where closed forms are available, and report experiments on composite nonlinear channels.
@@ II-A Notation
$J(U)\equiv\mathbb{E}_{\bm{u}\sim p_{U}(\cdot)}\!\left[\|\nabla_{\bm{u}}\log p_{U}(\bm{u})\|^{2}\right]$ (5)
@@ II-C AWGN channel and score function
$Y_{t}=X+Z_{t},\quad Z_{t}\sim\mathcal{N}(\mathbf{0},t\mathbf{I}_{n}),\quad t>0,$ (6)
where $t$ represents the noise variance. The matrix $\mathbf{I}_{n}$ is the $n\times n$ identity matrix. Let $p_{Y_{t}}$ denote the marginal density of $Y_{t}$. The score function of $p_{Y_{t}}$ is defined by
@@
where $\|\cdot\|$ denotes the Euclidean norm. Under standard regularity conditions, the following Tweedie identity [10, 11, 12] holds for $t>0$:
@@ II-D I–MMSE relation
$Y_{\gamma}=\sqrt{\gamma}X+N,\quad N\sim\mathcal{N}(\mathbf{0},\mathbf{I}_{n}),\quad N\perp X,$ (10)
@@
The de Bruijn identity was discussed in [15, 16], which is proved by using the heat partial differential equation.
@@ II-F Denoising score matching (DSM) method
$\bm{v}=\bm{w}+\sigma\bm{\epsilon},\quad\bm{w}\sim p,\quad\bm{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I}_{n}).$ (15)
@@ III Estimating Mutual Information in AWGN Channels
$Y_{t}=X+Z_{t},\quad Z_{t}\sim\mathcal{N}(\mathbf{0},t\mathbf{I}_{n}),\quad t>0,$ (19)
@@ III-C Estimation of Fisher Information
$\widehat{J}_{t}\triangleq\frac{1}{N}\sum_{i=1}^{N}\|s_{Y_{t}}(\bm{y}_{i})\|^{2}$ (29)
@@ III-D Gaussian input case for validation
In this subsection we provide closed-form expressions for the Gaussian-input case $X\sim\mathcal{N}(\mathbf{0},P\,\mathbf{I}_{n})$ in the AWGN channel family $Y_{t}=X+Z_{t}$ with $Z_{t}\sim\mathcal{N}(\mathbf{0},t\,\mathbf{I}_{n})$ where $P>0$ represents the signal power. It is well known [2] that $I(X;Y_{t})=(n/2)\log(1+P/t)$; see also [13] for the relation to MMSE.
@@
Let $X\sim\mathcal{N}(\mathbf{0},P\,\mathbf{I}_{n})$ and $Z_{t}\sim\mathcal{N}(\mathbf{0},t\,\mathbf{I}_{n})$, independent. Then, for all $t>0$, we have the following closed forms:
@@
Sketch of proof: .
From the Gaussian model $Y_{t}=X+Z_{t}$, it is known that $I(X;Y_{t})=(n/2)\log(1+P/t)$ holds. Moreover $Y_{t}\sim\mathcal{N}(\mathbf{0},(P+t)\mathbf{I}_{n})$, hence we have
@@ III-E Linear Gaussian channel case
$Y_{t}=AX+Z_{t},\qquad Z_{t}\sim\mathcal{N}(\mathbf{0},t\mathbf{I}_{n}),$ (34)
where $A\in\mathbb{R}^{n\times n}$ is a square, invertible matrix and $X\sim\mathcal{N}(\mathbf{0},P\mathbf{I}_{n})$. This setting preserves the Gaussian structure while introducing nontrivial correlation through the linear mixing due to $A$.
@@
Let $X\sim\mathcal{N}(\mathbf{0},P\mathbf{I}_{n})$, $A\in\mathbb{R}^{n\times n}$ nonsingular, and $Z_{t}\sim\mathcal{N}(\mathbf{0},t\mathbf{I}_{n})$, independent. Then for all $t>0$:
$I(X;Y_{t})$ $=\tfrac{1}{2}\log\det\!\Bigl(I_{n}+\tfrac{P}{t}AA^{\top}\Bigr),$ (35)
$\mathrm{mmse}(t)$ $=\mathrm{tr}\!\left(\Bigl(P^{-1}I_{n}+\tfrac{1}{t}A^{\top}A\Bigr)^{-1}\right),$ (36)
$J(Y_{t})$ $=\mathrm{tr}\!\Bigl((PAA^{\top}+tI_{n})^{-1}\Bigr).$ (37)
Sketch of proof:.
Since $AX$ is Gaussian with covariance $PAA^{\top}$, we have $Y_{t}\sim\mathcal{N}(\mathbf{0},PAA^{\top}+tI_{n})$. The mutual information (35) follows from the determinant formula for Gaussian differential entropy [2, 14]. The Fisher information (37) is obtained directly from the Gaussian score function $s_{Y_{t}}(\bm{y})=-(PAA^{\top}+tI_{n})^{-1}\bm{y}$. The MMSE formula (36) comes from the linear MMSE estimator of $X$ given $Y_{t}$. ∎
@@ IV-A Score-based Estimation Method for MI
In the above algorithm, we consider two alternatives for the DSM learning:
@@ V-B Score-based estimation method for nonlinear channels
In the following discussin, we assume that the evalution of $f$ is computationally feasible. Namely, sampling from $p_{W}$ is feasible. Some parts of Algorithm 1 should be replaced by the following steps:
@@ VI-A Gaussian input case for validation
To validate the proposed score-based estimation method, we first consider the Gaussian input case where exact closed-form expressions are available. Specifically, we consider the Gaussian input case $X\sim\mathcal{N}(0,PI_{n})$ with the AWGN channel model $Y_{t}=X+\sqrt{t}Z$, where $Z\sim\mathcal{N}(0,I_{n})$.
@@ VI-C Linear Gaussian channel case
$Y_{t}={\bm{A}}X+Z_{t},\qquad Z_{t}\sim\mathcal{N}(\mathbf{0},t\mathbf{I}_{n}),$ (48)
where ${\bm{A}}\in\mathbb{R}^{n\times n}$ is a random orthogonal matrix and $X\sim\mathcal{N}(\mathbf{0},P\mathbf{I}_{n})$ with $P=1.0$, $n=4$.

### v3

@@ Abstract
We present a numerical method to evaluate mutual information (MI) in nonlinear Gaussian noise channels by using denoising score matching (DSM) learning for estimating the score function of channel output. Via de Bruijn’s identity, Fisher information estimated from the learned score function yields accurate estimates of MI through a Fisher integral representation for a variety of priors and channel nonlinearities. In this work, we propose a comprehensive theoretical foundation for the Score-to-Fisher bridge methodology, along with practical guidelines for its implementation. We also conduct extensive validation experiments, comparing our approach with closed-form solutions and a kernel density estimation baseline. The results of our numerical experiments demonstrate that the proposed method is both practical and efficient for MI estimation in nonlinear Gaussian noise channels. Additionally, we discuss the theoretical connections between our score-based framework and thermodynamic concepts, such as partition function estimation and optimal transport.
@@ I Introduction
$Y_{t}=f(X)+Z_{t},\qquad Z_{t}\sim\mathcal{N}(\bm{0},t\bm{I}_{n}),$ (1)
@@
so that the mutual information $I(W;Y_{t})$ can be obtained by integrating the Fisher information $J(Y_{t})$.
@@
In this paper, we propose a score-based estimator of MI [^1] A software implementation to reproduce the numerical results in this paper is available at https://github.com/wadayama/score-to-fisher-mi. tailored to the channel represented by $Y_{t}=f(X)+Z_{t}$. The method learns the score of channel output, or the marginal density $p_{Y_{t}}$, using DSM learning and then obtain an estimate of the Fisher information $\widehat{J}(Y_{t}).$ From this, we obtain an estimate of the mutual information via an Fisher integral representation (4). Crucially, the approach never requires access to the posterior $p_{X|Y}(\cdot|\cdot)$. All we need is possibility of forward channel simulation, i.e., we assume that generation of samples $\bm{x}\sim p_{X}(\cdot)$, evaluations of $\bm{w}=f(\bm{x})$, and Gaussian noise addition are computationally feasible.
The main contributions of this paper are summarized as follows: (1)Score-only MI estimation for nonlinear front-ends: we formalize a pipeline: DSM learning of $p_{Y_{t}}$ $\Rightarrow$ Fisher estimation $\Rightarrow$ Fisher integral representation—that applies to arbitrary deterministic $f$ as long as $p_{Y_{t}}$ can be easily sampled. (2)Practical recipe: we detail implementation choices that make the estimator stable and reproducible: logarithmic grids in $t$, tail corrections for the MI integral. (3)Validation and use-cases: we verify the estimator on Gaussian inputs where closed forms are available, and report experiments on composite nonlinear channels.
@@
In modern score-based modeling, the score function model serves as a tractable surrogate for otherwise intractable densities. Our score-to-fisher bridge leverages this view to turn mutual information into a one-dimensional integral of score-based functionals along Gaussian smoothing, enabling accurate numerical MI estimation without high-dimensional density estimation.
@@ II-A Notation
$J(U)\equiv\mathbb{E}_{\bm{u}\sim p_{U}(\cdot)}\!\left[\|\nabla_{\bm{u}}\log p_{U}(\bm{u})\|^{2}\right],$ (5)
@@ II-B1 Estimation of Mutual Information
II-B2 Theoretical Foundations and Thermodynamic Connections
The integral representation used in this work is deeply rooted in the interplay between information theory and thermodynamics, particularly through the heat equation. This connection dates back to Stam [15], who derived the de Bruijn identity to prove the entropy power inequality, explicitly treating noise addition as a diffusion process. This line of inquiry was further developed by Brown [16] and Barron [17], who established the entropic central limit theorem, proving that the Fisher information and entropy monotonically converge to those of a Gaussian distribution during diffusion. In a modern context, Guo, Shamai, and Verdú [13] unified these concepts through the I-MMSE relationship, revealing that the derivative of mutual information is equivalent to the minimum mean-square error (and thus Fisher information) in Gaussian channels. Our proposed Score-to-Fisher bridge can be viewed as a data-driven operationalization of these classical theoretical frameworks: while Stam and Barron utilized these relations for proofs of inequalities, we leverage them for the exact numerical estimation of information quantities via learned score functions.
II-B3 Diffusion Process and Fisher Information
Recently, score-based representations have found applications beyond generative modeling, extending into information-theoretic analysis. For instance, Premkumar [23] introduced “neural entropy” to quantify the information stored in diffusion models, relating the total entropy production along a diffusion path to a time integral of a Fisher-divergence functional (i.e., an expected squared difference of score functions). While their framework is motivated by a non-equilibrium thermodynamic view of diffusion generative models (and the resulting quantity depends on the chosen forward process), it shares a fundamental mathematical connection with our approach: both convert global information measures into one-dimensional integrals of score-based functionals, through diffusion/heat-equation identities (de Bruijn’s identity in our additive Gaussian-noise setting).
@@ II-C AWGN channel and score function
$Y_{t}=X+Z_{t},\quad Z_{t}\sim\mathcal{N}(\bm{0},t\bm{I}_{n}),\quad t>0,$ (6)
where $t$ represents the noise variance. The matrix $\bm{I}_{n}$ is the $n\times n$ identity matrix. Let $p_{Y_{t}}$ denote the marginal density of $Y_{t}$. The score function of $p_{Y_{t}}$ is defined by
@@
where $\|\cdot\|$ denotes the Euclidean norm. Under standard regularity conditions, the following Robbins-Tweedie identity [10, 11, 12] holds for $t>0$:
@@ II-D I–MMSE relation
$Y_{\gamma}=\sqrt{\gamma}X+N,\quad N\sim\mathcal{N}(\bm{0},\bm{I}_{n}),\quad N\perp X,$ (10)
@@
The de Bruijn identity was discussed in [15, 17], which is proved by using the heat partial differential equation. Intuitively, diffusion acts as a smoothing process that destroys information. As illustrated in Fig. 1, this process is characterized by the decay of Fisher information $J(Y_{t})$ (which measures the ‘sharpness’ or structural complexity of the distribution) and the simultaneous increase in entropy $h(Y_{t})$. Our method computes the total information content by integrating this Fisher information profile.
Figure 1: Visualization of the heat diffusion process for a 1D multimodal distribution (mixture of Gaussians). As time $t$ progresses, the distribution $p(x,t)$ becomes smoother, leading to an increase in differential entropy $h(X_{t})$ and a rapid decay in Fisher information $J(X_{t})$. The Score-to-Fisher Bridge method leverages this relationship via de Bruijn’s identity: the mutual information is essentially accumulated Fisher-information gap relative to the Gaussian baseline (the complexity of the distribution) over the diffusion time.
@@ II-F Denoising score matching (DSM) method
$\bm{v}=\bm{w}+\sigma\bm{\epsilon},\quad\bm{w}\sim p,\quad\bm{\epsilon}\sim\mathcal{N}(\bm{0},\bm{I}_{n}).$ (15)
@@ III Estimating Mutual Information in AWGN Channels
$Y_{t}=X+Z_{t},\quad Z_{t}\sim\mathcal{N}(\bm{0},t\bm{I}_{n}),\quad t>0,$ (19)
@@ III-B Fisher Integral representations of mutual information
Remark 1 (Connection to GEXIT function).
The integrand in our estimation formula corresponds to the channel GEXIT function $g(t)$. Although typically defined via conditional entropy, it satisfies the following identities in the AWGN setting:
$g(t)\equiv\frac{d}{dt}h(X|Y_{t})=-\frac{d}{dt}I(X;Y_{t})=\frac{n}{2t}-\frac{1}{2}J(Y_{t}).$ (26)
By the area theorem, the integral of $g(t)$ characterizes the mutual information, thus linking our score-based Fisher integral to the GEXIT analysis commonly used in coding theory.
@@ III-C Estimation of Fisher Information
$\widehat{J}_{t}\equiv\frac{1}{N}\sum_{i=1}^{N}\|s_{\bm{\theta}}(\bm{y}_{i})\|^{2}$ (30)
@@ III-D Gaussian input case for validation
In this subsection we provide closed-form expressions for the Gaussian-input case $X\sim\mathcal{N}(\bm{0},P\,\bm{I}_{n})$ in the AWGN channel family $Y_{t}=X+Z_{t}$ with $Z_{t}\sim\mathcal{N}(\bm{0},t\,\bm{I}_{n})$ where $P>0$ represents the signal power. It is well known [2] that $I(X;Y_{t})=(n/2)\log(1+P/t)$; see also [13] for the relation to MMSE.
@@
Let $X\sim\mathcal{N}(\bm{0},P\,\bm{I}_{n})$ and $Z_{t}\sim\mathcal{N}(\bm{0},t\,\bm{I}_{n})$, independent. Then, for all $t>0$, we have the following closed forms:
@@
Proof:
From the Gaussian model $Y_{t}=X+Z_{t}$, it is known that $I(X;Y_{t})=(n/2)\log(1+P/t)$ holds. Moreover $Y_{t}\sim\mathcal{N}(\bm{0},(P+t)\bm{I}_{n})$, hence we have
@@ III-E Linear Gaussian channel case
$Y_{t}=\bm{A}X+Z_{t},\qquad Z_{t}\sim\mathcal{N}(\bm{0},t\bm{I}_{n}),$ (35)
where $\bm{A}\in\mathbb{R}^{n\times n}$ is a square, invertible matrix and $X\sim\mathcal{N}(\bm{0},P\bm{I}_{n})$. This setting preserves the Gaussian structure while introducing nontrivial correlation through the linear mixing due to $A$.
@@
Let $X\sim\mathcal{N}(\bm{0},P\bm{I}_{n})$, $\bm{A}\in\mathbb{R}^{n\times n}$ nonsingular, and $Z_{t}\sim\mathcal{N}(\bm{0},t\bm{I}_{n})$, independent. Then for all $t>0$:
$I(X;Y_{t})$ $=\tfrac{1}{2}\log\det\!\Bigl(I_{n}+\tfrac{P}{t}\bm{A}\bm{A}^{\top}\Bigr),$ (36)
$\mathrm{mmse}(t)$ $=\mathrm{tr}\!\left(\Bigl(P^{-1}I_{n}+\tfrac{1}{t}\bm{A}^{\top}\bm{A}\Bigr)^{-1}\right),$ (37)
$J(Y_{t})$ $=\mathrm{tr}\!\Bigl((P\bm{A}\bm{A}^{\top}+t\bm{I}_{n})^{-1}\Bigr).$ (38)
Proof:
Since $AX$ is Gaussian with covariance $P\bm{A}\bm{A}^{\top}$, we have $Y_{t}\sim\mathcal{N}(\bm{0},P\bm{A}\bm{A}^{\top}+t\bm{I}_{n})$. The mutual information (36) follows from the determinant formula for Gaussian differential entropy [2, 14]. The Fisher information (38) is obtained directly from the Gaussian score function $s_{Y_{t}}(\bm{y})=-(P\bm{A}\bm{A}^{\top}+t\bm{I}_{n})^{-1}\bm{y}$. The MMSE formula (37) comes from the linear MMSE estimator of $X$ given $Y_{t}$. ∎
@@ IV-A Score-based Estimation Method for MI
It should be noted that if the probability density $p_{Y_{t}}$ of the channel output $Y_{t}$ is known, and its score function $s_{Y_{t}}(y)=\nabla_{y}\log p_{Y_{t}}(y)$ is available in closed form (e.g., in the case of an identity function $f$ with a Gaussian Mixture Model (GMM) input $X$), then Step 3 (DSM learning) becomes unnecessary. In such cases, one can directly use this analytical score function in Step 4 to perform the Monte Carlo estimation of the Fisher information. This approach eliminates the score approximation error while retaining the advantages over conventional methods in terms of computational cost and high-dimensional scalability. However, the primary focus of this paper is on more general nonlinear channels where such an analytical solution is intractable or unavailable.
in Algorithm 1, we consider two alternatives for the DSM learning:
@@ IV-B Log-domain trapezoid integration
IV-C Stein Calibration
In practice, the learned score function $s_{\theta}(\mathbf{y})$ may exhibit a scaling bias. To mitigate this, we apply Stein calibration to the score estimates. Leveraging Stein’s identity, we compute a scalar correction factor $c$ such that the calibrated score $\tilde{s}_{\theta}(\mathbf{y})=cs_{\theta}(\mathbf{y})$ satisfies the empirical moment constraint $\mathbb{E}[\mathbf{Y}_{t}^{\top}\tilde{s}_{\theta}(\mathbf{Y}_{t})]=-n$, assuming centered data. The estimated Fisher information is then computed using this calibrated score.
IV-D Alternative Score Estimation via Velocity Fields
While we employed DSM to estimate the score function, the SFB framework is inherently estimator-agnostic. Recent generative paradigms, such as flow matching [24] and rectified flow [25], learn a velocity field that drives the probability flow, typically along straight-line trajectories. Since the score function can be mathematically derived from the learned velocity field (given the specific scaling factor), these methods offer a promising alternative route for score estimation. Leveraging such techniques could potentially improve the numerical stability of MI estimation, particularly in regimes where DSM struggles with curvature in the diffusion path. See Appendix G for more details.
@@ V-B Score-based estimation method for nonlinear channels
In the following discussion, we assume that the evaluation of $f$ is computationally feasible. Namely, sampling from $p_{W}$ is feasible. Some parts of Algorithm 1 should be replaced by the following steps:
@@
V-C Error Analysis and Convergence Rates
This section provides a qualitative analysis of the estimation error for the proposed estimator $\hat{I}$. The total estimation error can be decomposed into four distinct components:
$|\hat{I}-I|\;\leq\;\varepsilon_{\mathrm{score}}+\varepsilon_{\mathrm{MC}}+\varepsilon_{\mathrm{quad}}+\varepsilon_{\mathrm{tail}}.$
Each component behaves as follows under standard regularity assumptions:
  1. 1.
    Score Approximation Error ($\varepsilon_{\mathrm{score}}$): This term arises from the imperfect learning of the score function $s_{\theta}\approx\nabla\log p_{t}$. It scales with the square root of the Fisher divergence loss. Improving the neural network capacity and training budget reduces this bias.
  2. 2.
    Monte Carlo Variance ($\varepsilon_{\mathrm{MC}}$): Since the Fisher information $J(Y_{t})$ is estimated via sample averaging, this stochastic error follows the standard Central Limit Theorem scaling of $O(N^{-1/2})$, where $N$ is the batch size per grid point.
  3. 3.
    Quadrature Error ($\varepsilon_{\mathrm{quad}}$): The discretization of the integral over $u=\log t$ using the trapezoidal rule introduces a deterministic error. On a uniform grid with $M$ points (step size $\Delta u\propto M^{-1}$), this error decays rapidly as $O(M^{-2})$.
  4. 4.
    Truncation Error ($\varepsilon_{\mathrm{tail}}$): Restricting the integration to $t_{\min},t_{\max}]$ introduces bias from the missing tails. The error typically decays as $O(t_{\max}^{-1})$ for the large-$t$ tail and polynomially (e.g., $O(t_{\min}^{\beta})$) for the small-$t$ tail.
To minimize the total error efficiently, one must balance these components. Since the quadrature error $\varepsilon_{\mathrm{quad}}\sim O(M^{-2})$ decays much faster than the Monte Carlo error $\varepsilon_{\mathrm{MC}}\sim O(N^{-1/2})$, a relatively coarse grid $M$ is often sufficient compared to the sample size $N$. Theoretically, balancing the rates suggests a scaling of $M\asymp N^{1/5}$. In practice, we recommend fixing a sufficiently wide integration range $t_{\min},t_{\max}]$ to suppress $\varepsilon_{\mathrm{tail}}$, and then prioritizing a large batch size $N$ to mitigate the dominant Monte Carlo variance.
@@ VI-A Gaussian input case for validation
To validate the proposed score-based estimation method, we first consider the Gaussian input case where exact closed-form expressions are available. Specifically, we consider the Gaussian input case $X\sim\mathcal{N}(\bm{0},P\bm{I}_{n})$ with the AWGN channel model $Y_{t}=X+\sqrt{t}Z$, where $Z\sim\mathcal{N}(\bm{0},\bm{I}_{n})$.
@@ VI-C Linear Gaussian channel case
$Y_{t}={\bm{A}}X+Z_{t},\qquad Z_{t}\sim\mathcal{N}(\bm{0},t\bm{I}_{n}),$ (49)
where ${\bm{A}}\in\mathbb{R}^{n\times n}$ is a random orthogonal matrix and $X\sim\mathcal{N}(\bm{0},P\bm{I}_{n})$ with $P=1.0$, $n=4$.